OpenStudy (anonymous):

Solve -3x2 - 4x - 4 = 0

OpenStudy (anonymous):

is it -3 times 2 or are you saying -3x?

OpenStudy (anonymous):

it is \[-3x ^{2}\]

OpenStudy (anonymous):

ok give me one min ill solve the problem

OpenStudy (anonymous):

ok thanks.all i know of it is it is going to be a complex number solution

OpenStudy (anonymous):

ok so what I got was 9x but im really not sure ill try to show you the steps one sec

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wait wait wait im wrong

OpenStudy (anonymous):

ill redo do it sorry

OpenStudy (anonymous):

it ok man it confuses the crap outa me.

OpenStudy (anonymous):

ok lets try to do it together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-3x is squared making it 9x right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok so then 9x - 4x is 5x

OpenStudy (anonymous):

then you are left with 4 or 4x

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

what grade is this for?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry i was away from keyboard.it is 10th grade algebra 2

OpenStudy (anonymous):

oh wow haha im only in 8th grade

OpenStudy (anonymous):

well sorry that i couldn't be much help

OpenStudy (anonymous):

lmfao XD

OpenStudy (anonymous):

sorry :(

OpenStudy (anonymous):

it ok dude

OpenStudy (anonymous):

use the quadratic formula: \[\frac{ 4 \pm \sqrt{(-4)^2-4(-3)(-4)} }{ 2(-3) }\] reduces to: \[\frac{ 4 \pm \sqrt{-32} }{ -6 }\] since you cannot take the square root of a negative you insert i. \[x =\frac{ 4 + \sqrt{32}i }{ -6 }\] AND \[x =\frac{ 4 - \sqrt{32}i }{ -6 }\]

OpenStudy (anonymous):

thanks piglet9

OpenStudy (anonymous):

No Problem =D